What charge is stored in a 100 μF capacitor when 120 V is applied to it?...

  1. What charge is stored in a 100 μF capacitor when 120 V is applied to it?
  2. Find the charge stores when 5.50 V is applied to an 8.00 pF capacitor.
  3. Calculate the voltage applied to a 2.00 μF capacitor when it holds 3.10 μC charge.
Answer #1

Answer:

1)

Given, C ( Capacitance)=100microfarad= 100×10^-6 Farad

Voltage, V=120v

Q= C V

Q(charge) =100×10^-6 ×120

Q=12×10^-3 C

2) Given, voltage, V=5.50 v

C(capacitance) =8 pF =8×10^-12 F

Q(charge) =44 ×10^-12 C

3)

Given, C, capacitance=2 microfarad =2×10^-6 F

Charge (Q) =3.10 Microcoulomb= 3.10×10^-6C

Q=C V

V= Q/C =3.10 ×10^-6 /2 ×10^-6

V=1.55 volts